# REOrdering Patches Improves Vision Models

### v1

@@ 6 Learning to Order Patches
Our findings in Section 4 reveal that patch ordering can significantly affect the performance of long-sequence vision models. This suggests the potential value of discovering a data- and model-specific ordering to improve task performance. Unfortunately, learning a discrete permutation poses a unique challenge. For an image with $N$ patches, there are $N!$ possible orderings. For $N=196$, there are more than $10^{365}$ options. Searching this combinatorial space exhaustively is infeasible. A naïve approach would require evaluating each permutation’s classification loss on every training example, which is computationally intractable and incompatible with gradient-based learning. We instead treat the selection of patch orderings as a stochastic policy learning problem, where the policy outputs a distribution over permutations. This allows us to sample permutations during training and optimize the policy against the downstream classification loss with reinforcement learning.
@@ 6.1 Learning the Permutation Policy via REINFORCE
3:$\pi\leftarrow\operatorname*{argsort}_{\operatorname{desc}}(z+\tau g)$triangleright Gumbel-top-$k$ permutation
4:$x_{\pi}^{(b)}\leftarrow\textsc{Permute}(x^{(b)},\pi)\quad\forall b$
5:$\hat{y}^{(b)}\leftarrow f_{\theta}\bigl{(}x_{\pi}^{(b)}\bigr{)}$
6:$\mathcal{L}_{\operatorname{CE}}\leftarrow-\frac{1}{B}\sum_{b=1}^{B}\log\hat{y}^{(b)}_{\,y^{(b)}}$
7:$r\leftarrow-\mathcal{L}_{\operatorname{CE}}$triangleright Reward
8:$b\leftarrow\beta b+(1-\beta)r$
9:$A\leftarrow r-b$triangleright Advantage
10:$\mathcal{L}_{\operatorname{policy}}\leftarrow-A\,\log P(\pi\mid z)$
11:$\mathcal{L}_{\operatorname{total}}\leftarrow\mathcal{L}_{\operatorname{CE}}+\mathcal{L}_{\operatorname{policy}}$
@@ 6.2 The Plackett-Luce Policy for Patch Orderings
To parameterize the permutation distribution, we use the Plackett-Luce (PL) model luceIndividualChoiceBehavior1959 ; plackettAnalysisPermutations1975 . This model defines a distribution over permutations based on a learned logit vector $z\in\mathbb{R}^{n}$ associated with the image patches. A permutation is sampled sequentially: at each step, a yet-unplaced patch is selected according to a softmax over the remaining logits:
@@
The logit vector $z$ is initialized as a linear ramp from 0 to $-1$, then permuted according to the information-theoretic prior described in Section 5. This gives the model a sensible starting point that reflects structural cues discovered during unsupervised analysis, while still allowing gradients to adapt the ordering throughout training.
@@
Figure 4: The logits of the Plackett-Luce model, and therefore the permutation order, changes over the course of training. Longformer is initialized with column and row-major patch ordering and optimized with REOrder. The image is of the class “keyboard.” We track two patches over the course of the policy curriculum: a keyboard key (light red arrow) and an irrelevant orange beak (dark red arrow). As the policy learns to order patches, we see the patches related to the target class move to the end of the ordering.
@@ 6.3 Curriculum for Policy Learning
We adopt a three-stage curriculum that cleanly separates representation learning from policy learning. For the first $N$ epochs the classifier is trained with the canonical row-major patch order. This gives the vision transformer a stable starting point before any permutation noise is introduced. Beginning at epoch $N$ the Plackett-Luce policy is activated and trained with REINFORCE for $M$ epochs. Sampling uses the Gumbel-top-$k$ trick with a temperature schedule $\tau_{t}$ that climbs to a peak value before decaying to zero. While $\tau_{t}>0$ the permutation remains stochastic, encouraging exploration and allowing the policy to discover beneficial orderings. Once the temperature hits zero at epoch $N+M$, sampling collapses to the single maximum-likelihood permutation $\hat{\pi}=\operatorname*{argsort}(z)$. The policy gradients vanish, freezing the permutation, and the remaining epochs let the backbone finish optimizing with its now-determinate input order. The evolution of the policy is visualized in Figure 4, demonstrating how salient patches for the target class of “keyboard” move to the end of the sequence to maximize classification accuracy.
@@ 6.4 Effectiveness of Learned Patch Ordering
The Plackett-Luce policy introduced by REOrder is a simple drop-in addition to model training, and as visualized in Figure 5, in almost all cases, improves performance over their respective base patch order runs. Mamba observes gains of $2.20\%~{}(\pm 0.22\%)$ on average across different patch orderings for IN1K, and $9.32\%~{}(\pm 0.22\%)$ on FMoW. For ImageNet, the Hilbert curve ordering for Mamba is improved by $3.01\%~{}(\pm 0.23\%)$ while for Functional Map of the World, the diagonal ordering is improved by $13.35\%~{}(\pm 0.21\%)$. Transformer-XL demonstrates modest gains with REOrder. On IN1K, Transformer-XL accuracy improves by an average of $0.70\%~{}(\pm 0.22\%)$ but sees significant gains with the Hilbert curve ($1.50\%~{}(\pm 0.21\%)$) and spiral ($1.09\%~{}(\pm 0.21\%)$) patch orderings. On FMoW, REOrder improves the best-performing column-major order by an additional $1.10\%~{}(\pm 0.21\%)$, demonstrating that simply using a basic patch ordering alone may not be sufficient to get best performance. Longformer is unable to improve its accuracy on either dataset. However, since Longformer was the model that was initially also the least susceptible to changes in patch ordering due to its near-full approximation of self-attention, it is unsurprising that the use of REOrder does not achieve any noticeable performance gains.
@@
In summary, REOrder’s approach of learning a patch ordering proves to be a broadly effective strategy for boosting classification performance with long-sequence models, consistently elevating results even beyond the strongest baseline patch orderings for nearly all tested models and datasets.
@@ Appendix A Discussion and Limitations
A random baseline for a learned permutation baseline involves sampling a random ordering of patches for every batch during training. We set up this baseline with Transformer-XL on the ImageNet-1K dataset. This model achieved a maximum top-1 accuracy of 39.07% which is 15.25% worse than the worst patch ordering tested for T-XL on IN1K (spiral). While a random baseline should be run for every model and dataset combination, the drastically lower performance of the random baseline gives us confidence in the veracity of our results.
@@
Theorem 1 (Permutation equivariance of self-attention).
Let $X\in\mathbb{R}^{n\times d}$ and let $P\in\{0,1\}^{n\times n}$ be any permutation matrix. With $\operatorname{Attn}(\cdot)$ defined in Eq. (1), we have
$\operatorname{Attn}(PX)=P\,\operatorname{Attn}(X).$
@@
$S(X)=\operatorname{softmax}\mathopen{}\left(\frac{{(XW_{q})(XW_{k})^{\!\top}}}{{\sqrt{d}}}\right)\in\mathbb{R}^{n\times n}.$
Conventionally, softmax for $\operatorname{Attn}$ is applied row-wise over $(XW_{q})(XW_{k})^{\!\top}$ so that the ensuing multiplication by the value matrix serves as a normalized weighted sum over value columns. Since a permutation merely re-orders rows and columns, it satisfies the conjugation property
$\operatorname{softmax}(PMP^{\!\top})=P\,\operatorname{softmax}(M)\,P^{\!\top}$ (5)
for any square matrix $M$ and any permutation matrix $P$.
@@
$\operatorname{Attn}(PX)$ $=\operatorname{softmax}\!\Bigl{(}\tfrac{(PXW_{q})(PXW_{k})^{\!\top}}{\sqrt{d}}\Bigr{)}\,PXW_{v}$
$=\operatorname{softmax}\!\Bigl{(}\tfrac{P(XW_{q})(XW_{k})^{\!\top}P^{\!\top}}{\sqrt{d}}\Bigr{)}\,PXW_{v}$
=(5)${=}P\,S(X)\,P^{\!\top}\,PXW_{v}$
$=P\,S(X)\,XW_{v}$ $(P^{\!\top}P=I)$
=(1)${=}P\,\operatorname{Attn}(X).$ ) .
Hence self-attention is permutation equivariant. ∎
@@ Appendix D Additional Model Training Details
Each training run required 100 epochs for completion. ImageNet experiments were run on 8$\times$ 80GB A100 GPUs, while Functional Map of the World experiments were run on 4$\times$ 40GB A100 GPUs. Transformer-XL, ViT and Mamba, and the Plackett-Luce policy were compiled with TorchDynamo using the Inductor backend. The Longformer encoder provided by HuggingFace was unable to be compiled. For ImageNet runs, Vision Transformer runs took ~$8.5$ hours to run, Transformer-XL runs took ~$10$ hours, Longformer runs took ~$12$ hours, and Mamba runs took ~$19$ hours. For Functional Map of the World, Vision Transformer runs took ~$17.5$ hours, Transformer-XL runs took ~$21$ hours, Longformer runs took ~$26$ hours, and Mamba runs took ~$31$ hours. All runs were executed on dedicated datacenters accessed remotely.
@@
D.1 Results Without Flips
@@ F.6 Snake Order
Let $s_{k}=r_{k}+c_{k}$ be the anti-diagonal index for the $k$-th patch $\pi(k)=(r_{k},c_{k})$. The sequence of patches $\pi(k)$ is generated by iterating $s$ from 00 to $H+W-2$. For each $s$:
@@ Appendix G Attention Patterns for Tested Models
Figure 6: Token-level attention coverage across different model architectures. Each grid cell represents a patch in a 224$\times$224 input image split into 49 non-overlapping 32$\times$32 patches, plus a [CLS] token. Numbers indicate how many tokens attend to each patch. ViT and Mamba exhibit full attention to all patches. In contrast, Transformer-XL’s causal attention and Longformer’s local attention restrict the number of tokens that can attend to each patch, leading to a strong asymmetry and localized attention, respectively.
To produce Figure 6, we visualized how many tokens attend to each patch in a 224$\times$224 image divided into 49 patches (plus a [CLS] token), using the following methods for each model.

### v2

@@ 6 Learning to Order Patches
Our findings in Section˜4 reveal that patch ordering can significantly affect the performance of long-sequence vision models. This suggests the potential value of discovering a data- and model-specific ordering to improve task performance. Unfortunately, learning a discrete permutation poses a unique challenge. For an image with $N$ patches, there are $N!$ possible orderings. For $N=196$, there are more than $10^{365}$ options. Searching this combinatorial space exhaustively is infeasible. A naïve approach would require evaluating each permutation’s classification loss on every training example, which is computationally intractable and incompatible with gradient-based learning. We instead treat the selection of patch orderings as a stochastic policy learning problem, where the policy outputs a distribution over permutations. This allows us to sample permutations during training and optimize the policy against the downstream classification loss with reinforcement learning.
@@ 6.1 Learning the Permutation Policy via REINFORCE
[…]
@@ 6.2 The Plackett-Luce Policy for Patch Orderings
To parameterize the permutation distribution, we use the Plackett-Luce (PL) model luceIndividualChoiceBehavior1959 ; plackettAnalysisPermutations1975 . This model defines a distribution over permutations based on a learned logit vector $z\in\mathbb{R}^{n}$ associated with the image patches. Each patch gets a single parameter, so a $\times224$ image with patch size $\times16$ results in a model with $196$ parameters. This is negligible added overhead for training. A permutation is sampled sequentially: at each step, a yet-unplaced patch is selected according to a softmax over the remaining logits:
@@
The logit vector $z$ is initialized as a linear ramp from 0 to 1-1, then permuted according to the information-theoretic prior described in Section˜5. This gives the model a sensible starting point that reflects structural cues discovered during unsupervised analysis, while still allowing gradients to adapt the ordering throughout training.
@@
Figure 4: The logits of the Plackett-Luce model, and therefore the permutation order, changes over the course of training. Longformer is initialized with column and row-major patch ordering and optimized with REOrder on ImageNet-1K. The image is of the class “keyboard.” We track two patches over the course of the policy curriculum: a keyboard key (light red arrow) and an irrelevant orange beak (dark red arrow). As the policy learns to order patches, we see the patches move toward the back of the sequence (i.e., are back-loaded) reflecting the dataset’s center bias.
@@ 6.3 Curriculum for Policy Learning
We adopt a three-stage curriculum that cleanly separates representation learning from policy learning. For the first $N$ epochs the classifier is trained with the canonical row-major patch order. This gives the vision transformer a stable starting point before any permutation noise is introduced. Beginning at epoch $N$ the Plackett-Luce policy is activated and trained with REINFORCE for $M$ epochs. Sampling uses the Gumbel-top-$k$ trick with a temperature schedule $\tau_{t}$ that climbs to a peak value before decaying to zero. While $\tau_{t}>0$ the permutation remains stochastic, encouraging exploration and allowing the policy to discover beneficial orderings. Once the temperature hits zero at epoch $N+M$, sampling collapses to the single maximum-likelihood permutation $\hat{\pi}=\operatorname*{argsort}(z)$. The policy gradients vanish, freezing the permutation, and the remaining epochs let the backbone finish optimizing with its now-determinate input order.
For our experiments. the curriculum was set with $N=15$ and $M=30$. $\tau_{t}$ follows a triangle pattern where it increases linearly from $\tau=0.0$ at epoch $N$ to $M/2$ with $\tau=0.2$. It then decreases linearly from epoch $M/2$ to $M$ ending at $\tau=0$. The optimizer for the policy was AdamW($\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$, weight decay $0.03$) with a learning rate of $1\times 10^{-4}$. The evolution of the policy over training is visualized in Figure˜4, demonstrating how salient patches for the target class of “keyboard” move to the end of the sequence to maximize classification accuracy.
@@ 6.4 Effectiveness of Learned Patch Ordering
The Plackett-Luce policy introduced by REOrder is a simple drop-in addition to model training, and as visualized in Figure˜5, in almost all cases, improves performance over their respective base patch order runs. Mamba observes gains of $2.20\%~(\pm 0.22$ percentage points)) on average across different patch orderings for IN1K, and $9.32\%~(\pm 0.22$ percentage points)) on FMoW. For ImageNet, the Hilbert curve ordering for Mamba is improved by $3.01\%~(\pm 0.23$ percentage points)) while for Functional Map of the World, the diagonal ordering is improved by $13.35\%~(\pm 0.21$ percentage points)). Transformer-XL demonstrates modest gains with REOrder. On IN1K, Transformer-XL accuracy improves by an average of $0.70\%~(\pm 0.22$ percentage points)) but sees significant gains with the Hilbert curve ($1.50\%~(\pm 0.21$ percentage points))) and spiral ($1.09\%~(\pm 0.21$ percentage points))) patch orderings. On FMoW, REOrder improves the best-performing column-major order by an additional $1.10\%~(\pm 0.21$ percentage points)), demonstrating that simply using a basic patch ordering alone may not be sufficient to get best performance. Longformer is unable to improve its accuracy on either dataset. However, since Longformer was the model that was initially also the least susceptible to changes in patch ordering due to its near-full approximation of self-attention, it is unsurprising that the use of REOrder does not achieve any noticeable performance gains.
@@
To verify that these gains are not artifacts of training dynamics or limited training duration, we conducted two complementary robustness analyses (see Appendix D.2 and D.1). The first examines the role of policy-driven exploration by comparing static and dynamic random permutation strategies. The second extends Transformer-XL training from 100 to 300 epochs to test stability under longer optimization schedules. Together, these experiments confirm that performance improvements arise from the reinforcement learning process itself and persist under extended training.
In summary, REOrder’s approach of learning a patch ordering proves to be a broadly effective strategy for boosting classification performance with long-sequence models, consistently elevating results even beyond the strongest baseline patch orderings for nearly all tested models and datasets. We outline current limitations of our study and potential future directions in  Appendix˜A. We also provide a thorough interpretability analysis in  Appendix˜H.
@@ Appendix A Discussion and Limitations
Directional Asymmetry and ARM Interpretation.
While our work primarily focuses on improving the permutation sensitivity of sequence-based vision models through REOrder, it is important to contextualize the architectural behavior of ARM. Although ARM’s four-directional scanning mechanism is symmetric in principle, our experiments show that its performance varies across different input orderings due to the anisotropic statistics of natural images and the model’s fixed directional composition. In particular, ARM achieves higher accuracy with a row-major ordering than with a column-major one, reflecting the stronger horizontal correlations inherent in natural scenes. This empirical asymmetry persists despite theoretical symmetry between the two configurations, revealing how data bias and architectural inductive bias jointly shape model behavior. Moreover, a Hilbert-curve ordering, which maximizes spatial locality, performs significantly worse because it disrupts ARM’s directional continuity. ARM’s recurrence dynamics are optimized for specific scanning patterns rather than spatial adjacency alone.
REOrder addresses these mismatches by learning an adaptive permutation of the input sequence that jointly benefits all four directional scans. When initialized from a row-major order, it improves validation top-1 accuracy to $58.25\%$, and from a column-major order to $58.58\%$ on ImageNet-1K. Even from a suboptimal Hilbert initialization, REOrder recovers much of the lost performance, reaching $53.42\%$. These results indicate that the learned permutations do not merely refine an existing scan but instead adaptively reorganize the patch sequence to better align the input statistics with ARM’s directional processing. This demonstrates that REOrder can serve as a practical mechanism for reconciling architectural symmetry with anisotropic visual data.
@@
A random baseline for a learned permutation baseline involves sampling a random ordering of patches for every batch during training. We set up this baseline with Transformer-XL on the ImageNet-1K dataset. This model achieved a maximum top-1 accuracy of 39.0739.07% which is $15.25$ percentage points worse than the worst patch ordering tested for T-XL on IN1K (spiral). While a random baseline should be run for every model and dataset combination, the drastically lower performance of the random baseline gives us confidence in the veracity of our results.
@@
Dynamic Image Policy.
A key limitation of the present work is that REOrder learns a single global ordering shared across all samples within a dataset. While this design enables clean isolation of ordering effects and ensures stable optimization, it constrains the model’s flexibility in adapting to image-specific structures. A promising future direction is to extend REOrder toward dynamic, per-image ordering, where a lightweight policy network predicts ordering logits conditioned on patch embeddings. Such an approach would allow the ordering to adapt to spatial and semantic content, potentially capturing intra-dataset variability while preserving the benefits of learned sequencing. Investigating the trade-offs between global stability and dynamic adaptability represents an important next step for our future work.
Long Sequence Training.
To assess scalability, we initiated experiments on ultra–high-resolution images ($5888{\times}5888$), a setting where standard ViT models exceed GPU memory limits. These experiments are ongoing, and we plan to include results in future work exploring how REOrder behaves in extended sequence regimes.
@@
Theorem 1 (Permutation equivariance of self‑attention).
Let $\mathbf{X}\in\mathbb{R}^{n\times d}$ and let $\mathbf{P}\in\{0,1\}^{n\times n}$ be any permutation matrix. With $\operatorname{Attn}(\cdot)$ defined in Eq. (1), we have
$\operatorname{Attn}(\mathbf{P}\mathbf{X})=\mathbf{P}\,\operatorname{Attn}(\mathbf{X}).$
@@
$S(\mathbf{X})=\operatorname{softmax}\mathopen{}\left(\frac{{(\mathbf{X}\mathbf{W}_{q})(\mathbf{X}\mathbf{W}_{k})^{\!\top}}}{{\sqrt{d}}}\right)\in\mathbb{R}^{n\times n}.$
Conventionally, softmax for $\operatorname{Attn}$ is applied row-wise over $(\mathbf{X}\mathbf{W}_{q})(\mathbf{X}\mathbf{W}_{k})^{\!\top}$ so that the ensuing multiplication by the value matrix serves as a normalized weighted sum over value columns. Since a permutation merely re‑orders rows and columns, it satisfies the conjugation property
$\operatorname{softmax}(\mathbf{P}\mathbf{M}\mathbf{P}^{\!\top})=\mathbf{P}\,\operatorname{softmax}(\mathbf{M})\,\mathbf{P}^{\!\top}$ (5)
for any square matrix $\mathbf{M}$ and any permutation matrix $\mathbf{M}$.
@@
$\operatorname{Attn}(\mathbf{P}\mathbf{X})$ $=\operatorname{softmax}\!\Bigl(\tfrac{(\mathbf{P}\mathbf{X}\mathbf{W}_{q})(\mathbf{P}\mathbf{X}\mathbf{W}_{k})^{\!\top}}{\sqrt{d}}\Bigr)\,\mathbf{P}\mathbf{X}\mathbf{W}_{v}$
$=\operatorname{softmax}\!\Bigl(\tfrac{\mathbf{P}(\mathbf{X}\mathbf{W}_{q})(\mathbf{X}\mathbf{W}_{k})^{\!\top}\mathbf{P}^{\!\top}}{\sqrt{d}}\Bigr)\,\mathbf{P}\mathbf{X}\mathbf{W}_{v}$
[…]
$=\mathbf{P}\,S(\mathbf{X})\,\mathbf{X}\mathbf{W}_{v}$ $(\mathbf{P}^{\!\top}\mathbf{P}=\mathbf{I})$
[…]
Hence self‑attention is permutation equivariant. ∎
@@ Appendix D Additional Model Training Details
Each training run required 100 epochs for completion. ImageNet experiments were run on $\times 8$0GB A100 GPUs, while Functional Map of the World experiments were run on $\times 4$0GB A100 GPUs. Transformer-XL, ViT and Mamba, and the Plackett-Luce policy were compiled with TorchDynamo using the Inductor backend. The Longformer encoder provided by HuggingFace was unable to be compiled. For ImageNet runs, Vision Transformer runs took ~$8.5$ hours to run, Transformer-XL runs took ~$10$ hours, Longformer runs took ~$12$ hours, and Mamba runs took ~$19$ hours. For Functional Map of the World, Vision Transformer runs took ~$17.5$ hours, Transformer-XL runs took ~$21$ hours, Longformer runs took ~$26$ hours, and Mamba runs took ~$31$ hours. All runs were executed on dedicated datacenters accessed remotely.
@@
D.1 Increased Training Time
We continued training our Transformer-XL Row and Transformer-XL REOrder Row models from 100 epochs to 300 epochs to test the robustness of our conclusions with longer training. The models were trained with the same recipe and learning curriculum. The gains from REOrder are maintained and even increased after 300 epochs. At 100 epochs (Table 3), Transformer-XL Row achieves $60.80\%$ Top-1 Validation accuracy whereas REOrder improves on this by $0.6$ percentage points to $61.40\%$. Extending training to 300 epochs, Transformer-XL Row achieves $61.20\%$, a $1.19$ percentage points increase over 100 epochs; however, Transformer-XL REOrder Row, increases performance to $63.34\%$ which is a $2.14$ percentage points increase over the Transformer-XL Row model at the same epoch. These findings show that REOrder continues to provide consistent gains under extended optimization, supporting its robustness beyond the initial training regime.
Table 3: REOrder widens the gap between standard models when training is extended. When training is extended from 100 epochs to 300 epochs, REOrder still shows significant Top-1 Validation Accuracy gains over training without REOrder.
Model Dataset Patch Order RL Top-1 Validation Accuracy Gain at 300 Epochs
T-XL IN1K Row - 60.80 +1.19pp1.19\,$p$
T-XL IN1K Row 61.40 +1.94pp1.94\,$p$
D.2 Random and Static Permutation Results
To better understand the effects of the REOrder learning process on the performance of the models we performed three additional runs with Transformer-XL on ImageNet-1k:
  1. 1.
    Using a fixed random permutation for the entire training
  2. 2.
    Sampling a random permutation for every batch through all of training
  3. 3.
    Using the final learn permutation of our best performing REOrder Transformer-XL model for all of training
The results in Table 4 indicate that simply picking random patch orderings is not sufficient (and therefore, any one random ordering will also perform poorly). The structured exploration provided by REOrder is necessary to improve performance. A novel outcome of this exploration is that simply using a learned patch ordering is not sufficient for gains in performance; the guided exploration process introduced by REOrder is necessary for improved performance.
Table 4: Top-1 validation accuracy on ImageNet-1K using Transformer-XL under four alternative permutation schemes. Each variant controls how patch sequences are presented during training. Only REOrder’s reinforcement-learning curriculum (second row) adaptively explores the ordering space, leading to a large performance gain over all static or unguided random strategies.
Permutation Strategy RL Top-1 Accuracy
Row Major - 60.80
REOrder Row-major initialization 61.40
One random permutation for all epochs - 53.10
Every batch random - 50.23
Best learned ordering - 53.07
D.3 Results Without Flips
@@ F.6 Snake Order
Let $s_{k}=r_{k}+c_{k}$ be the anti-diagonal index for the $k$-th patch $\pi(k)=(r_{k},c_{k})$. The sequence of patches $\pi(k)$ is generated by iterating $s$ from 0 to H+W2H+W-2. For each $s$:
@@ Appendix G Attention Patterns for Tested Models
Figure 6: Token-level attention coverage across different model architectures. Each grid cell represents a patch in a $\times224$ input image split into 49 non-overlapping $\times32$ patches, plus a [CLS] token. Numbers indicate how many tokens attend to each patch. ViT and Mamba exhibit full attention to all patches. In contrast, Transformer-XL’s causal attention and Longformer’s local attention restrict the number of tokens that can attend to each patch, leading to a strong asymmetry and localized attention, respectively.
To produce Figure 6, we visualized how many tokens attend to each patch in a $\times224$ image divided into 49 patches (plus a [CLS] token), using the following methods for each model.